# The Gauss map of a projective foliation

### v2

@@ 2. Basic properties, Results, and Examples
so that
$\mathcal{G}_{\mathcal{F}}^{-1}\circ\mathcal{G}_{\mathcal{F}}(z)(z)=-z_{0}\sum_{j=1}^{n}z_{j}^{2}\,(z_{0},z_{1},...,z_{n})=-\ell\,q.\,z$
As a consequence $\mathcal{G}_{\mathcal{F}}$ contracts the hypersurfaces $(\ell=0)$ and $(q=0)$, which are leaves of $\mathcal{F}$. The hyperplane $(x_{0}=0)$ is contracted to the point $[1:0:...:0]$ and the quadric $(q=0)$ to the quadric of codimension two $(y_{0}=0)\cap(q(y)=0)$. Note that this case, in dimension three, is one of the homaloids, that we will study in example 6.
@@
Note that this foliation is in the irreducible component $\mathcal{L}(1,...,1;n)$ ($n+1$ ones) of the space of codimension one foliations on $\mathbb{P}^{n}$ of degree $n-1$. The Gauss map in this case is
@@
which is birational and whose inverse has a similar expression, this is a self-dual foliation. The map $\mathcal{G}_{\mathcal{F}}$ contracts the union of the hyperplanes $(z_{j}=0)$, $0\leq j\leq n$.
Remark 2.7.
Here we will see that for any degree $d\geq 1$ there are associated foliations on $\mathbb{P}^{3}$ whose associated Gauss map is birational.
@@
$F_{k}(x,y,z,t)=\frac{g}{y^{k+2}}=\frac{(xz+yt)y^{k}+x^{k+2}}{y^{k+2}}$
It is defined by the 1-form
@@
$=y[(k+2)x^{k+1}+y^{k}z]dx-[(k+2)x^{k+2}-2xy^{k}z-y^{k+1}t]dy+xy^{k+1}\,dz+y^{k+2}\,dt,$
and the Gauss map is   $\mathcal{G}_{\mathcal{F}_{k}}[x:y:z:t]=$
$=[y[(k+2)x^{k+1}+y^{k}z]:-[(k+2)x^{k+2}-2xy^{k}z-y^{k+1}t]:xy^{k+1}:y^{k+2}]\,.$
In this example we have $Sing_{2}(\mathcal{F})=(x=y=0):=\ell$ and that the hyperplane $\Sigma=(y=0)$ is $\mathcal{F}$-invariant and contracted to a point by $\mathcal{G}_{\mathcal{F}_{k}}$.
In the affine coordinates $(y=1)\simeq\mathbb{C}^{3}$ the foliation has the first integral
$f_{k}(x,z,t)=F_{k}(x,1,z,t)=t+xz+x^{k+2}\,.$
Consider the hyperplane $H=(t=\alpha x+\beta z+\gamma)$. It can be checked directly that $f_{k}|_{H}$ has a unique singularity of Morse type. This implies that $\mathcal{G}_{\mathcal{F}_{k}}$ is birational. We would like to remark that if $k=1$ then $\mathcal{F}_{1}\in\mathcal{L}(3,1;3)\subset\mathbb{F}\text{ol}(2;3)$. For a generic member $\mathcal{F}\in\mathcal{L}(3,1;3)$ it can be shown that $tdg(\mathcal{G}_{\mathcal{F}})=3$, so that $\mathcal{G}_{\mathcal{F}}$ is not birational (see example 9).
If $k\geq 2$ then $\mathcal{F}_{k}\in\mathcal{L}(k+2,1;3)\subset\mathbb{F}\text{ol}(k+1;3)$. For a generic member of $\mathcal{F}\in\mathcal{L}(k+1,3;3)$ then $Sing_{2}(\mathcal{F})$ is a plane curve of degree $k+2$, but in the case of $\mathcal{F}_{k}$ then $Sing_{2}(\mathcal{F}_{k})$ degenerates in a straight line of multiplicity $k+2$.
@@
Magic example.   Here is an example for which the associated foliation has no special transversal structure and contracts hypersurfaces that are not invariant for the foliation. Let $\omega=A(u,v)du+B(u,v)dv$, where $A$ and $B$ are non zero polynomials. We consider the pull-back map $\Phi(x,y,z,t)=(x/t,y/z)$. Note that
$\Phi^{*}(\omega)=A(x/t,y/z)d\left(\frac{x}{t}\right)+B(x/t,y/z)d\left(\frac{y}{z}\right):=\theta$
We have
@@
We note that, in general, the original foliation in $\mathbb{P}^{2}$, has no invariant algebraic curve as in the Jouanolou foliation [27]. Sometimes can even have some dense leaves (see [28] and [29]).
With this method, we can obtain foliations on $\mathbb{P}^{3}$ with Gauss map of any desired degree $\geq 2$, without rational first integral.
@@
In example 4 all fibers of the pull-back map $[x:y:z:t]\dasharrow(x/t,y/z)$ are straight lines. A natural question is: what happens if the fibers are not plane curves?
@@
$\mathcal{G}_{\mathcal{F}}^{-1}[A:B:C:D]$ $:\displaystyle=[2\lambda ABC((\lambda-1)B-C):\lambda C((2-\lambda)A^{2}B+4B^{2}D+A^{2}C)$ :
$B(\lambda A^{2}+4\lambda BD)(C-\lambda B):4\lambda B^{2}C((\lambda-1)B-C)]$
The facts to be noted in this example are the following: The fibers of $F$ are conic plane curves, the hypersurface given by $\{yt+x^{2}-zt=0\}$ is not invariant by $\mathcal{F}$, however, it is contracted by the Gauss map $\mathcal{G}_{\mathcal{F}}$. Finally, we can also note that for $\lambda$ generic, the foliation $\mathcal{F}$ has no rational first integral.
@@
Before stating the next results let us recall some facts about a codimension one foliation $\mathcal{F}$ on $\mathbb{P}^{n}$, $n\geq 3$.
The singular set of $\mathcal{F}$ always has irreducible components of codimension two (cf [24]). We will denote as $Sing_{2}(\mathcal{F})$ the union of the irreducible components of codimension two of $Sing(\mathcal{F})$.
Let $V\subset Sing_{2}(\mathcal{F})$ be an irreducible component of codimension two. We define the transversal multiplicity of $V$ as follows: fix a smooth point $p\in V$ and a transverse 2-dimensional disk $\Sigma$ through $p$ such that $\Sigma$ is transverse to $\mathcal{F}$ outside $p$. Then the restriction $\mathcal{F}|_{\Sigma}$, is a foliation on $\Sigma$ with an isolated singularity at $p$. The multiplicity of $\mathcal{F}$ at $(\Sigma,p)$ is the Milnor’s number $\mu(\mathcal{F}|_{\Sigma},p)$ (see [5]).
We define the multiplicity of $\mathcal{F}$ at $p\in V$ as:
@@
It can be proved that the function $p\in V\mapsto\mu(\mathcal{F},V,p)\in\mathbb{N}$ is constant in a Zariski open and dense subset of $V$. This constant will be denoted as $\mu(\mathcal{F},V)$. It is called the transverse multiplicity of $\mathcal{F}$ at $V$.
Remark 2.8.
@@
For instance, when $\mathcal{F}$ is defined by a logarithmic form that is the case (see proposition 4).
@@
  • (a).
    $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j}).dg(V_{j})\leq d^{2}+d+1$. The equality is attained when $\mathcal{F}$ is a pull-back of a foliation on $\mathbb{P}^{2}$ by a linear map $\Psi\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{2}$ (see [9]).
  • (b).
    Let $m(\mathcal{F})=d^{2}+d+1-\sum_{j=1}^{k}\mu(\mathcal{F},V_{j}).dg(V_{j})$. Assume that $m(\mathcal{F})>0$ and that $\mathcal{G}_{\mathcal{F}}$ is dominant. Then $\mathcal{G}_{\mathcal{F}}\colon\mathbb{P}^{3}\dasharrow\breve{\mathbb{P}}^{3}$ has topological degree $m(\mathcal{F})$, that is, it is generically $m(\mathcal{F})$ to one. In particular, $\mathcal{G}_{\mathcal{F}}$ is birational if, and only if $\mathcal{G}_{\mathcal{F}}$ is dominant and $m(\mathcal{F})=1$.
Corollary 2.2.
@@
Corollary 2.3.
Let $\mathcal{F}$ be a foliation on $\mathbb{P}^{n}$ of degree $d$, $n\geq 4$. Let $V_{1},...,V_{k}$ be the irreducible components of $Sing_{2}(\mathcal{F})$. Assume that $\mathcal{G}_{\mathcal{F}}$ is dominant and that $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j})\,deg(V_{j})=d^{2}+d$. Then $\mathcal{G}_{\mathcal{F}}$ is birational.
Corollary 2.4.
Let $\mathcal{F}$ be a foliation on $\mathbb{P}^{n}$ of degree $d$, $n\geq 4$. Let $V_{1},...,V_{k}$ be the irreducible components of $Sing_{2}(\mathcal{F})$. Assume that $\mathcal{G}_{\mathcal{F}}$ is dominant and that $\sum_{j=1}^{k}deg(V_{j})=d^{2}+d$. Then $\mathcal{G}_{\mathcal{F}}$ is birational.
Remark 2.9.
We would like to note that the hypothesis of corollary 2.3  is sufficient for the Gauss map of a foliation on $\mathbb{P}^{n}$ to be birational, but is not necessary when $n\geq 4$. For instance, in the case of example 3, where $\omega=z_{0}z_{1}...z_{n}\sum_{j=0}^{n}\mu_{j}\frac{dz_{j}}{z_{j}}$, we have:
@@ 3. Since $deg(\mathcal{F}_{\omega})=n-1:=d$ we get
As the reader can check, we have $m=1$ if, and only if, $n=3$. Therefore if $n\geq 4$ the condition in corollary 2.3 is not necessary.
@@
All these examples can be verified using theorem 3. Let us verify case (3). In homogeneous coordinates, it has a rational first integral of the form $F(x,y,z,t)=\frac{y(x^{2}+yz)}{t^{3}}$ whose foliation $\mathcal{F}$ is defined by the form
@@
$=\frac{1}{y^{p+2}t^{q+2}}\left(x^{p}yt^{q+2}dx-x^{p+1}t^{q+2}dy+y^{p+2}z^{q}tdz-y^{p+2}z^{q+1}dt\right)$
@@
where the $M_{j^{\prime}s}$ are monomials of the same degree; $M_{j}(z)=\lambda_{j}\,z^{\sigma_{j}}$. As usually, we use the notations $\sigma=(\sigma(0),...,\sigma(n))$ and $z^{\sigma}=z_{0}^{\sigma(0)}.z_{1}^{\sigma(1)}...z_{n}^{\sigma(n)}$. We will assume that in (3) the monomials are not scalar multiples of one particular monomial: $[M_{0}:M_{1}:...:M_{n}]\neq[\lambda_{0}:\lambda_{1}:...:\lambda_{n}]\in\mathbb{P}^{n}$. A typical example in $\mathbb{P}^{3}$ is the one with $\omega=u^{p}du+v^{q}dv$ with $u=x/y$ and $v=z/t$.
@@
Let $\omega=\sum_{i=0}^{3}A_{i}(z)dz_{i}$ be a monomial integrable 1-form on $\mathbb{C}^{4}$, defining a monomial foliation $\mathcal{F}_{\omega}$ on $\mathbb{P}^{3}$ of degree $\geq 2$. If the Gauss map $\mathcal{G}_{\omega}$ of $\mathcal{F}_{\omega}$ is dominant then either $\mathcal{F}_{\omega}$ belongs to $\mathcal{L}(1,1,1,1;3)$; or it is a pull-back of a form on $\mathbb{C}^{2}$ like in example 4, for degree $\geq 3$. In particular, $\mathcal{G}_{\omega}\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{3}$ is birational.
@@
Here is an application of theorem 3. The exceptional foliation of degree two on $\mathbb{P}^{3}$, $\mathcal{E}\in\mathbb{F}\text{ol}(2;3)$, is defined in homogeneous coordinates by the form
@@
In homogeneous coordinates, it has a rational first integral of the form $F=f^{2}/g^{3}$, where $deg(f)=3$ and $deg(g)=2$. It is called exceptional because the logarithmic form $\frac{dF}{F}=2\frac{df}{f}-3\frac{dg}{g}$ has $t$ as divisor of zeroes. We note that $f(x,y,z,t)=zt^{2}-xyt+x^{3}/3$, $g(x,y,z,t)=yt-x^{2}/2$ (see [9]) and $fg\frac{dF}{F}=t\omega$ and $cod(Sing(\omega))=2$, so that $\omega$ defines $\mathcal{E}$ in homogeneous coordinates. We would like to observe that it was proved in [9] that the irreducible component of $\mathbb{F}\text{ol}(2,3)$ that contains the exceptional foliation is the closure of the orbit of $\mathcal{E}$ by the natural action of $Aut(\mathbb{P}^{3})$. From now on this component will be denoted as $\mathcal{E}(2,3;3)$.
@@
It is proven also that $\mu(\mathcal{F},V_{j})=1$, $j=1,2,3$. Therefore, in this case we have $m(\mathcal{E})=7-(3+2+1)=1$ and $\mathcal{G}_{\mathcal{E}}$ is birational. The inverse $\mathcal{G}_{\mathcal{E}}^{-1}$ will be done in § 4.1.
This example will be studied in detail in § 4.1 and in § 4.2.
@@
We would like to observe that this example was originally introduced in [8]. It was proved in [12] that the irreducible component of $\mathbb{F}\text{ol}(3;4)$ that contains $\mathcal{E}^{\prime}$ is the closure of its orbit by the natural action of $Aut(\mathbb{P}^{4})$. From now on this component will be denoted as $\mathcal{E}(3,4;4)$. We define $\mathcal{E}(3,4;n)$, $n\geq 4$, as the set
@@
It can be proved that $\mathcal{E}(3,4;n)$ is an irreducible component of $\mathbb{F}\text{ol}(4;n)$ for all $n$. However, if $\mathcal{E}\in\mathcal{E}(3,4;n)$ and $n>4$ then $\mathcal{G}_{\mathcal{E}}$ is not dominant.
Remark 2.11.
In § 4.4 we will generalize these examples of exceptional foliations for foliations of degree $n-1$ on $\mathbb{P}^{n}$. As in the case of examples 7 and 8, these examples will have a rational first integral of the form $f^{n-1}/g^{n}$, where $deg(f)=n$ and $deg(g)=n-1$.

### v3

@@ 2. Basic properties, Results, and Examples
thus
$\mathcal{G}_{\mathcal{F}}^{-1}\circ\mathcal{G}_{\mathcal{F}}(z)=-z_{0}\sum_{j=1}^{n}z_{j}^{2}\,(z_{0},z_{1},...,z_{n})=-\ell\,q.\,z$
As a consequence $\mathcal{G}_{\mathcal{F}}$ contracts the hypersurfaces $(\ell=0)$ and $(q=0)$, which are leaves of $\mathcal{F}$. The hyperplane $(z_{0}=0)$ is contracted to the point [1:0::0][1:0:...:0] and the quadric $(q=0)$ to the quadric of codimension two $(y_{0}=0)\cap(q(y)=0)$. Note that this case, in dimension three, is one of the homaloids, that we will study in example 6.
@@
Note that this foliation is in the irreducible component $\mathcal{L}(1,...,1;n)$ ($n+1$ ones) of the space of codimension one foliations on $\mathbb{P}^{n}$ of degree n1n-1. The Gauss map in this case is
@@
which is birational and whose inverse has a similar expression; this is a self-dual foliation. The map $\mathcal{G}_{\mathcal{F}}$ contracts the union of the hyperplanes $(z_{j}=0)$, $0\leq j\leq n$.
Remark 7.
Here we will see that for any degree $d\geq 1$ there are foliations on $\mathbb{P}^{3}$ whose associated Gauss map is birational.
@@
$F_{k}(x,y,z,t)=\frac{g}{y^{k+2}}=\frac{(xz+yt)y^{k}+x^{k+2}}{y^{k+2}}.$
Then the foliation is defined by the 1-form
@@
$y[(k+2)x^{k+1}+y^{k}z]dx-[(k+2)x^{k+2}-2xy^{k}z-y^{k+1}t]dy+xy^{k+1}\,dz+y^{k+2}\,dt,$
and the Gauss map is   $\mathcal{G}_{\mathcal{F}_{k}}[x:y:z:t]=[A:B:C:D]$
$y[(k+2)x^{k+1}+y^{k}z]:-[(k+2)x^{k+2}-2xy^{k}z-y^{k+1}t]:xy^{k+1}:y^{k+2}].\,$
In this example we have that the singular set is $(x=y=0):=\ell$, the hyperplane $\Sigma=(y=0)$ is $\mathcal{F}$-invariant and this is contracted to a point by $\mathcal{G}_{\mathcal{F}_{k}}$. This Gauss map is birational with inverse $\mathcal{G}^{-1}_{\mathcal{F}_{k}}[A:B:C:D]$
$CD^{k+1}:D^{k+2}:AD^{k+1}-(k+2)C^{k+1}D:BD^{k+1}-2ACD^{k}+3(k+2)C^{k+2}].\,$
We would like to remark that if $k=1$ then $\mathcal{F}_{1}\in\mathcal{L}(3,1;3)\subset\mathbb{F}\text{ol}(2;3)$. For a generic member $\mathcal{F}\in\mathcal{L}(3,1;3)$ it can be shown that $tdg(\mathcal{G}_{\mathcal{F}})=4$, so that $\mathcal{G}_{\mathcal{F}}$ is not birational (see example 10).
If $k\geq 2$ then $\mathcal{F}_{k}\in\mathcal{L}(k+2,1;3)\subset\mathbb{F}\text{ol}(k+1;3)$. For a generic member $\mathcal{F}\in\mathcal{L}(k+1,3;3)$ the singular set is a plane curve of degree $k+2$, but in the case of $\mathcal{F}_{k}$, the singular set degenerates in a straight line of multiplicity $k+2$.
@@
Magic example.   Here is an example for which the associated foliation has no special transversal structure and contracts hypersurfaces that are not invariant for the foliation. Let $\omega=A(u,v)du+B(u,v)dv$, where $A$ and $B$ are non-zero polynomials. We consider the pull-back map $\Phi(x,y,z,t)=(x/t,y/z)$. Note that
$\Phi^{*}(\omega)=A(x/t,y/z)d\left(\frac{x}{t}\right)+B(x/t,y/z)d\left(\frac{y}{z}\right):=\theta.$
Then we have
@@
We note that, in general, the original foliation in $\mathbb{P}^{2}$, has no invariant algebraic curves as in the Jouanolou foliation [28] (see chapter 4 of [23]). Sometimes can even have some dense leaves (see [29] and [30]).
With this method, we can obtain foliations on $\mathbb{P}^{3}$ with birational Gauss map of any desired degree $\geq 2$, without rational first integral.
To prove this statement, it suffices to consider $\omega=f(u)du+dv$, where $f(u)$ is a polynomial of degree $s$. Applying the pull-back by $\Phi$ in the affine chart $t=1$, we obtain $z^{2}f(x)dx+zdy-ydz$, which defines a foliation of degree $s+2$.
@@
In example 4 all fibers of the pull-back map $x:y:z:t]\dasharrow(x/t,y/z)$ are straight lines. A natural question is: what happens if the fibers are not plane curves?
@@
$\mathcal{G}_{\mathcal{F}}^{-1}[A:B:C:D]$ $=[2\lambda ABC((\lambda-1)B-C):\lambda C((2-\lambda)A^{2}B+4B^{2}D+A^{2}C):$
$B(\lambda A^{2}+4\lambda BD)(C-\lambda B):4\lambda B^{2}C((\lambda-1)B-C)].$
The facts to be noted in this example are the following: The fibers of $F$ are conic plane curves, the hypersurface given by $\{yt+x^{2}-zt=0\}$ is not invariant by $\mathcal{F}$, however, it is contracted by the Gauss map $\mathcal{G}_{\mathcal{F}}$. Finally, we can also note that for $\lambda$ generic, the foliation $\mathcal{F}$ has no rational first integral, because the 1-form $(u-v)du+\lambda udv$ does not have it.
@@
Before stating the next results, let us recall some facts about a codimension one foliation $\mathcal{F}$ on $\mathbb{P}^{n}$.
The singular set of $\mathcal{F}$ always has irreducible components of codimension two (see for example [26]). We will denote as $Sing_{2}(\mathcal{F})$ the union of the irreducible components of codimension two of $Sing(\mathcal{F})$.
If $\mathcal{F}$ is a foliation on $\mathbb{P}^{2}$, the Milnor number $\mu(\mathcal{F},p)$, of $\mathcal{F}$ at a point $p$ is defined as the dimension of the local ring at $p$ modulo the ideal generated by the local vector field defining the foliation near $p$ (see [6]).
Let $\mathcal{F}\in\mathbb{F}ol(d;n)$, where $n\geq 3$, and let $V\subset Sing_{2}(\mathcal{F})$ be an irreducible component of codimension two. Fix a smooth point $p\in V$ and a transverse 2-dimensional disk $\Sigma$ through $p$ such that $\Sigma$ is transverse to $\mathcal{F}$ outside $p$. Since the restriction $\mathcal{F}|_{\Sigma}$, is a foliation on $\Sigma$ with an isolated singularity at $p$ we can define the Milnor number $\mu(\mathcal{F}|_{\Sigma},p)$. It is called the multiplicity of $\mathcal{F}$ at $(\Sigma,p)$.
Definition 3.
We define the transversal multiplicity of $V$ as follows. Firstly we called:
@@
multiplicity of $\mathcal{F}$ at $p\in V$. It can be proved that the function $p\in V\mapsto\mu(\mathcal{F},V,p)\in\mathbb{N}$ is constant in a Zariski open and dense subset of $V$. This constant will be denoted as $\mu(\mathcal{F},V)$, and it is called the transverse multiplicity of $\mathcal{F}$ at $V$.
Remark 8.
@@
For instance, when $\mathcal{F}$ is defined by a generic logarithmic form that is the case (see proposition 4).
@@
  • (a)
    $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j}).deg(V_{j})\leq d^{2}+d+1$. The equality is attained when $\mathcal{F}$ is a pull-back of a foliation on $\mathbb{P}^{2}$ by a linear map $\Psi\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{2}$ (see [10] for a complete description of these foliations).
  • (b)
    Let $m(\mathcal{F})=d^{2}+d+1-\sum_{j=1}^{k}\mu(\mathcal{F},V_{j}).deg(V_{j})$, then $\mathcal{G}_{\mathcal{F}}\colon\mathbb{P}^{3}\dasharrow\breve{\mathbb{P}}^{3}$ has topological degree $m(\mathcal{F})$. In particular, $\mathcal{G}_{\mathcal{F}}$ is dominant if and only if, $m(\mathcal{F})>0$; and $\mathcal{G}_{\mathcal{F}}$ is birational if and only if, $m(\mathcal{F})=1$.
Corollary 2.
@@
The following three corollaries of the previous Theorem are valid in all dimensions, and give us sufficient conditions on the singular set of the foliation, to have birational Gauss map. The first and third are similar except for a small simplification of the formula described. The second one is for restrictions of foliations.
Corollary 3.
Let $\mathcal{F}$ be a foliation on $\mathbb{P}^{n}$ of degree $d$, $n\geq 4$. Let $V_{1},...,V_{k}$ be the irreducible components of $Sing_{2}(\mathcal{F})$. If $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j})\,deg(V_{j})=d^{2}+d$, then $\mathcal{G}_{\mathcal{F}}$ is birational.
Corollary 4.
Let $\mathcal{F}$ be a foliation on $\mathbb{P}^{n}$ of degree $d$, $n\geq 3$. Let $V_{1},...,V_{k}$ be the irreducible components of $Sing_{2}(\mathcal{F})$. Assume that $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j})\,deg(V_{j})=d^{2}+d$. Let $3\leq m\leq n$, and suppose that $\mathbb{P}^{m}$ is linearly embedded in $\mathbb{P}^{n}$, then the Gauss map of the restriction $\mathcal{F}_{|\mathbb{P}^{m}}$ is birational.
Corollary 5.
Let $\mathcal{F}$ be a foliation on $\mathbb{P}^{n}$ of degree $d$, $n\geq 4$. Let $V_{1},...,V_{k}$ be the irreducible components of $Sing_{2}(\mathcal{F})$. If $\sum_{j=1}^{k}deg(V_{j})=d^{2}+d$, then $\mathcal{G}_{\mathcal{F}}$ is birational.
Remark 9.
We would like to note that the hypothesis of corollary 3  is sufficient for the Gauss map of a foliation on $\mathbb{P}^{n}$ to be birational, but is not necessary when $n\geq 4$. For instance, in the case of example 3, where $\omega=z_{0}z_{1}...z_{n}\sum_{j=0}^{n}\mu_{j}\frac{dz_{j}}{z_{j}}$, we have:
@@ 3. Since $deg(\mathcal{F}_{\omega})=n-1:=d$ we get
As the reader can check, we have $m=1$ if, and only if, $n=3$. Therefore if $n\geq 4$ the condition in corollary 3 is not necessary.
@@
All these examples can be verified using Theorem 3. Let us verify case (3). In homogeneous coordinates, it has a rational first integral of the form $F(x,y,z,t)=\frac{y(x^{2}+yz)}{t^{3}}$, whose foliation $\mathcal{F}$ is defined by the form
@@
$\frac{1}{y^{p+2}t^{q+2}}\left(x^{p}yt^{q+2}dx-x^{p+1}t^{q+2}dy+y^{p+2}z^{q}tdz-y^{p+2}z^{q+1}dt\right)$
@@
where the $M_{j^{\prime}s}$ are monomials of the same degree; $M_{j}(z)=\lambda_{j}\,z^{\sigma_{j}}$. As usually, we use the notations $\sigma=(\sigma(0),...,\sigma(n))$ and $z^{\sigma}=z_{0}^{\sigma(0)}.z_{1}^{\sigma(1)}...z_{n}^{\sigma(n)}$. We will assume that in (3) the monomials are not scalar multiples of one particular monomial: $M_{0}:M_{1}:...:M_{n}]\neq[\lambda_{0}:\lambda_{1}:...:\lambda_{n}]\in\mathbb{P}^{n}$. A typical example in $\mathbb{P}^{3}$ is the one with $\omega=u^{p}du+v^{q}dv$ with $u=x/y$ and $v=z/t$.
@@
Let $\omega=\sum_{i=1}^{4}A_{i}(z)dz_{i}$ be a monomial integrable 1-form on $\mathbb{C}^{4}$, defining a monomial foliation $\mathcal{F}_{\omega}$ on $\mathbb{P}^{3}$ of degree $\geq 2$. If the Gauss map $\mathcal{G}_{\omega}$ of $\mathcal{F}_{\omega}$ is dominant then $\mathcal{F}_{\omega}$ belongs to $\mathcal{L}(1,1,1,1;3)$; or it is a pull-back of a form on $\mathbb{C}^{2}$ like in example 4, for degree $\geq 3$. In particular, $\mathcal{G}_{\omega}\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{3}$ is birational.
@@
Here is another application of theorem 3. The exceptional foliation of degree two on $\mathbb{P}^{3}$, $\mathcal{E}\in\mathbb{F}\text{ol}(2;3)$, is defined in homogeneous coordinates by the form
@@
In homogeneous coordinates, it has a rational first integral of the form $F=f^{2}/g^{3}$, where $deg(f)=3$ and $deg(g)=2$. It is called exceptional because the logarithmic form $\frac{dF}{F}=2\frac{df}{f}-3\frac{dg}{g}$ has $t$ as divisor of zeroes. We note that $f(x,y,z,t)=zt^{2}-xyt+x^{3}/3$, $g(x,y,z,t)=yt-x^{2}/2$ (see [10]), $fg\frac{dF}{F}=t\omega$ and $cod(Sing(\omega))=2$, so that $\omega$ defines $\mathcal{E}$ in homogeneous coordinates. We would like to observe that it was proved in [10] that the irreducible component of $\mathbb{F}\text{ol}(2,3)$ that contains the exceptional foliation is the closure of the orbit of $\mathcal{E}$ by the natural action of $Aut(\mathbb{P}^{3})$. From now on this component will be denoted as $\mathcal{E}(2,3;3)$.
@@
It is verified also that $\mu(\mathcal{F},V_{j})=1$, $j=1,2,3$. Therefore, in this case we have $m(\mathcal{E})=7-(3+2+1)=1$ and $\mathcal{G}_{\mathcal{E}}$ is birational. This example will be studied in detail in § 4.1 and in § 4.2. For example, the inverse $\mathcal{G}_{\mathcal{E}}^{-1}$ will be done in § 4.1.
@@
We would like to observe that this example was originally introduced in [9]. It was proved in [14] that the irreducible component of $\mathbb{F}\text{ol}(3;4)$ that contains $\mathcal{E}^{\prime}$ is the closure of its orbit by the natural action of $Aut(\mathbb{P}^{4})$. From now on this component will be denoted as $\mathcal{E}(3,4;4)$.
We define $\mathcal{E}(3,4;n)$, $n\geq 4$, as the set
@@
It can be proved that $\mathcal{E}(3,4;n)$ is an irreducible component of $\mathbb{F}\text{ol}(4;n)$ for all $n$ (see [14]). However, if $\mathcal{E}\in\mathcal{E}(3,4;n)$ and $n>4$ then $\mathcal{G}_{\mathcal{E}}$ is not dominant.
Remark 11.
In § 4.4 we will generalize these examples of exceptional foliations for foliations of degree n1n-1 on $\mathbb{P}^{n}$. As in the case of examples 7 and 8, these examples will have a rational first integral of the form $f^{n-1}/g^{n}$, where $deg(f)=n$ and deg(g)=n1deg(g)=n-1.